trying to find the slope-intercept form that passes thru these given points (-3,-3) (1,9)
Answer by lwsshak3(11628)

trying to find the slope-intercept form that passes thru these given points (-3,-3) (1,9)
***
slope-intercept form of equation for a line: y=mx+b, m=slope, b=y-intercept
for given problem:
slope,m=∆y/∆x=(9-(-3))/(1-(-3))=(9+3)/(1+3)=12/4=3
equation: y=3x+b
solve for b using coordinates from one of the given points,(1,9), on the line
9=3*1+b
b=6
equation: y=3x+6